What are two numbers that when you multiply equal 54 and when you add equal negative 21?
Answer by mananth(16946)

xy=54
x= 54/y
x+y=-21
substitute the value of x
54/y+y=-21
(54+y^2)/y=-21
(54+y^2)=-21y
y^2+21y+54=0
y^2+18y+3y+54=0
(y+18)(y+3)=0
y = -18 OR -3
x= -3 OR -18